# Apollo: An Lightweight Multilingual Medical LLM towards Democratizing Medical AI to 6B People

### v2

@@ Abstract
Figure 1: Linguistic areas covered by Apollo that supports English, Chinese, French, Spanish, Arabic, and Hindi.
@@ 1 Introduction
Rationale of Multilinguality in Medical LLMs The rationale of multilingual adaption in medical LLMs (Cox & Maryns, 2021; Pecina et al., 2014) could be twofold. Firstly, non-native English-speaking doctors often engage in bilingual learning through their native language and English from the outset, naturally introducing multilingual challenges in the learning process (Markó et al., 2006). Secondly, to better serve local communities, especially in countries and regions with scarce medical resources, medical aid based on local languages often achieves higher communication efficiency and acceptance (Brindley et al., 2014; Albrecht et al., 2013). Meanwhile, local medical knowledge can complement mainstream medical knowledge, fostering mutual benefits to accelerate medical development (Klayman, 1985; Yuan et al., 2016). Our pilot study in Sec. 2 also reveals that joint training of multiple languages enhances performance in the medical LLMs, indicating a beneficial complementarity among languages.
@@
The Lite LLM: Apollo The resulted multilingual medical LLMs trained by ApolloCorpora are named Apollo, to commemorate the Greek deity associated with healing, disease, the Sun, and light; this symbolizes the democratization of medical LLMs to 6 billion people, illuminating global healthcare. We explore a new domain adaption method that rewrite the pre-training corpora into QA pairs using ChatGPT (Li et al., 2023c) and adaptively sample training data, resulting in a smoother transition compared with the typical paradigm with continued pretraining and instruction tuning. Apollo ranges from 2B to 7B parameters. The advantage of the relatively-small model scale includes potential use as draft models for speculative decoding (Leviathan et al., 2023) or as proxy models for ProxyTuning (Liu et al., 2024a). In particular, we apply proxy tuning on top of Apollo to larger general LLMs, enhancing its multilingual medical capabilities. This is achieved without the need to directly train the general model using sensitive medical corpora, thereby underscoring the practical significance of Apollo in terms of protecting the privacy of medical training data against centralized training methods.
Figure 2: Apollo Pipeline Overview
@@
Contributions The contributions of this paper are as follows. (1) We collect and organize a high-quality multilingual medical corpora ApolloCorpora with rich local characteristics; (2) We obtain a series of SOTA multilingual LLMs Apollo at various parameter scales (especially in relatively-small sizes) ; (3) By using proxy tuning, Apollo could significantly improve larger general LLMs without finetuning, providing a new way to mitigating the exposure of private medical training data to centralized training systems; (4) We introduce a multilingual medical evaluation XMedBench and conduct extensive benchmark of existing models.
@@ 2.2 The Pilot Study
Model English Chinese French Spanish Arabic Hindi Avg.
Base Model
Qwen-1.8B 32.91 40.07 22.12 27.43 23.71 8.82 25.84
Language Specific Models
Apollo-English 39.44 45.27 28.35 31.76 22.61 8.72 29.36
Apollo-Chinese 39.42 61.13 28.97 33.83 27.94 25.34 36.11
Apollo-French 30.94 32.71 21.81 27.00 24.54 1.74 23.12
Apollo-Spanish 33.84 43.81 27.41 35.39 28.40 23.88 32.12
Apollo-Arabic 36.40 44.27 3.74 15.73 15.90 3.03 19.85
Apollo-Hindi 25.18 3.45 18.38 19.69 1.00 25.53 15.54
Our Method
Apollo (weight average) 40.54 45.58 28.04 34.08 28.95 24.06 33.54
Table 1: The pilot study on monolingual training and multilingual training.
Experimental settings To investigate the above question, we design a experimental protocol including the following three groups:
  • The backbone LM. We use a lite multilingual LLM Qwen 1.8B (Bai et al., 2023) as the LM backbone. It is argued that the findings should be agnostic to the selection of LM backbones, the selection of LLM Qwen 1.8B is due to its popularity, performance and more the importantly multilingual support.
@@
  • Monolingual training. The LM backbone is further trained by individual language, resulted six language-specific LLM variants (i.e. , English, Chinese, French, Spanish, Arabic, and Hindi.). The training data used could be found in Sec. 3. Moreover, we average the weights of these six LLM variants and obtain a new model that does not need further training, denoted as ‘weight average’.
  • Multilingual training. We train the LM using a mixture of the corpora in these six languages, denoted as ‘multilingual training’.
The evaluation protocol will be detailed in Sec. 4, which shows the average accuracy among datasets in each language.
Findings. Tab. 1 highlights the effectiveness of our methods in leveraging multilingual data to enhance the performance of medical Large Language Models (LLMs). The language-specific models, each trained exclusively on data from one language, demonstrate varying degrees of improvement in their respective languages over the original LM, underscoring the value of language-specific training. However, these models show limitations outside their target languages, as seen in the relatively low scores in non-target languages, particularly in the Apollo-French-1.8B and Apollo-Hindi-1.8B models. Our method, which includes both the weight average and multilingual training, significantly outperforms language-specific models across all languages in terms of average performance (see the last column). This illustrates the substantial benefits of combining multilingual data for training medical LLMs, with marked improvements in understanding and generating medical information across a diverse set of languages.
Therefore, we conclude the finding as below.
In general, multilingual medical corpora benefits medical LLMs.
3 ApolloCorpora, Multilingual Medical Dataset
In this chapter, we first introduce the organizational philosophy of the data set, then introduce the structure of the data set and the data collection process, then introduce the further processing of the data set, and finally analyze the local medical characteristics reflected in the data set.
3.1 Philosophy
Our goal is to enhance the multilingual medical capabilities of large-scale pre-trained models at an affordable computational cost, provide the community with a lightweight reproducible solution, and promote further exploration by the scientific research community. Considering that the model pre-training was most likely trained on PubMed-related data sets, we did not follow the mainstream method and repeat training on all Pubmed data.
If the model that has undergone large-scale pre-training is regarded as a medical student who is about to enter college with certain basic knowledge, in order to obtain more professional capabilities, we only need to replicate the high-quality data learned on the growth path of medical students. Starting from this point, after extensive communication with doctors and medical students, we identified six high-quality medical data collections: medical books, medical encyclopedias, medical clinical guidelines, medical papers, medical examinations, and professional doctor-patient dialogues. At the same time, in order to simulate the data input in the growth process of medical students in addition to medical professional knowledge, we also sampled medical-related parts of the Internet and collected instruction following data for some general tasks (including mathematical tasks of code).
For the dimension of multilingual medical expertise, we insist on using only medical data sets entirely from local languages and do not translate any medical-related data. This is done out of the following two considerations. First, there are many related works that prove that medical translation is a very complex task that cannot be simply solved by translation software; second, the expression habits of different languages, effective drugs, and even culture and Taboo terms arising from faith need to come from the local community intact, so as to maximize communication efficiency and avoid conflicts.
Regarding the License issue, we only screen data sets with complete open source protocols during the collection process to ensure that the open source protocols are friendly while ensuring quality.
@@ 3.1 Philosophy
3.2 Dataset Taxonomy and Collection
Books
For English books, we use medical dictionary  [^2] https://www.nlm.nih.gov/research/umls/new_users/online_learning/LEX_001.html to filter the books in the Pile Dataset (Gao et al., 2020) and select books with medical words accounting for more than 4%, and finally obtain 2312 medical-related books. For Chinese books, we follow MedQA (Jin et al., 2020) to collect medical textbooks included in the five-year and eight-year medical student training programs in mainland China, and finally obtain 90 books.
Papers
For English papers, we sample the public data in PubMed and obtain 878,241 medical abstracts. For Chinese papers, we also screen a total of 177,261 abstracts of papers published by the Chinese Medical Association  [^3] https://www.yiigle.com/index. For French papers, we use the MORFITT (Labrak et al., 2023b) dataset and the scientific article portion of the CLEAR (Grabar & Cardon, 2018). For the Spanish paper, we use paper abstracts open sourced by the Mesinesp (Gasco et al., 2021).
Encyclopedias
For the English Encyclopedia, we also use the English Medical Dictionary to filter out 36107 medical-related wiki pages from dataset [^4] https://huggingface.co/datasets/wikipedia. For the French encyclopedia, we select the encyclopedia articles part of the CLEAR (Grabar & Cardon, 2018). For the Hindi encyclopedia, we choose the HHD corpus (Jain & Arora, 2018), which crawls descriptions of people, diseases, medical consumer products, and symptoms from Indian websites.
Doctor-Patient Dialogues
For Chinese, we directly use the HuatuoGPT dataset (Zhang et al., 2023) and the simplified data set in Huatuo_26M (Li et al., 2023b). For English, we construct a multi-turn conversation data set based on PMC-Patients (Zhao et al., 2022) using ChatGPT, Prompt is shown in the Appendix A.1. For Arabic, we extract high-quality questions and answers with both question and answer lengths greater than 128 from the largest Arabic healthcare question and answer dataset MAQA (Abdelhay & Mohammed, 2022).
Exams
For the Chinese exam, we collect training sets of CMB (Wang et al., 2023), CMExam (Liu et al., 2024b), and MedQA (Zhang et al., 2018). For the English exam, we collect the training sets of MedQA, Medmcqa (Pal et al., 2022) and Pubmedqa (Jin et al., 2019). For the Spanish and French exam, we select the training set of HEAD-QA (Vilares & Gómez-Rodríguez, 2019) and Frenchmcqa (Labrak et al., 2023a) separately.
Guidelines
For English Guidelines, we select data from three sub-items of NICE [^5] https://www.nice.org.uk/guidance, PubMed and SPOR [^6] https://sporevidencealliance.ca/key-activities/cpg-asset-map/cpg-database/ in the clinical guidelines introduced by Meditron (Chen et al., 2023b).
General Instruction Tuning
We use the translation (FreedomIntelligence, 2023a) and original data of Sharegpt [^7] https://github.com/lm-sys/FastChat and Alpaca (Taori et al., 2023). For Chinese, we additionally make use of data (FreedomIntelligence, 2023b) generated by GPT-4 based on WizardLM Method (Xu et al., 2023). For English, in addition to adding the WizardLM Dataset, we also add belebele (Bandarkar et al., 2023) to enhance multi-language reading comprehension capabilities, ai2_arc (Clark et al., 2018) to enhance abstract reasoning capabilities, Capybara (Daniele & Suphavadeeprasit, 2023) to enhance instruction following capabilities.
Web
For Chinese, we use the medical dictionary (Han et al., 2016) to filter out medical-related articles from the Wudao Dataset (Xue et al., 2022). For English, we use the English Medical Vocabulary [^8] https://www.nlm.nih.gov/research/umls/new_users/online_learning/LEX_001.html to filter out medical related articles in C4 Dataset (Raffel et al., 2019). For Spanish, we sampled 10% of CoWeSe Dataset (Carrino et al., 2021).
Math
For mathematical abilities, we choose MathInstruct (Yue et al., 2023), a composite dataset containing various mathematics-related tasks and problem formats.
Code
We choose Python-Alpaca (Vezora, 2023) and Leetcode-ZH-11k (krisfu, 2023) respectively to strengthen the ability to solve coding tasks in Chinese and English.
3.3 Data processing
Rewrite
If we continue to train plain text, it is likely to destroy the original output format and knowledge. We want to explore whether rewriting the original pre-training corpus into QA pairs in the context of continuing training can help increase its medical capabilities without destroying the original model’s capabilities. We use ChatGPT999gpt-3.5-turbo-16k-0613 to generate questions and answers for a certain paragraph. For paragraph interception, we divide it according to the basic semantic units in the data set, such as sections in books and guides, paragraphs in website data, single wiki entry and abstracts of papers. For basic semantic units that are too long, we comprehensively consider the knowledge expression density of the language and subdivide different languages into blocks of different lengths to ensure that the semantic information covered by a single paragraph does not exceed the amount of information that can be included in a question and answer pair. For Spanish, French, English and Hindi we use 2048, for Chinese we use 256 and for Arabic we use 128. Prompts for generating QA pairs are detailed in the Appendix A.2.
Data Leakage Checking
The issue of data leakage is a recent focus of the academic community, which largely determines whether the results of the paper are convincing. For knowledge embedding tasks, data leakage screening with different stringency often leads to different performance. We follow Med-PaLM2 (Singhal et al., 2023) and adapt a more stringent deletion strategy. Specifically, we define a data item as leaked data if the entire question or at least 64 consecutive characters overlap with the data item. Regarding the exam exercise data source, there were 580,645 exercises before screening, and 3,041 exercises were deleted, with a screening rate of 0.52%. For other data sources, since they are not exam questions, there is no difference before and after filtering.
Examples Chinese: Representatives of exogenous wind-cold and high fever include: (A) Yinqiao Powder, (B) Qingwen Baidu Decoction, (C) Dachengqi Decoctio. (D) Xiaobuhu Decoction Arabic: Patient: “Peace be upon you, I feel very intense friction in my body and head, redness in the skin or its color changing to pink, and I am almost cut off in my body. This all happens when I start making an effort, even if it is the smallest of things, even if my voice rises or I get angry. Thank you, and may God’s peace, mercy, and blessings be upon you.” Doctor: “Often this is sensitivity due to exertion and sweating, and they are signs of a disease called fever of the nile, or a rash or allergy to sweating that appears in the summer season in particular. It is preferable to use a cold shower and apply the pimples with a weakening ointment. If the area is small, apply Calcipotriene ointment to it, take an anti-histamine medication such as Polaramine.” Hindi: Disease: Epilepsy. Tulsi reduces many diseases like a panacea. A large amount of anti-oxidants are found in Tulsi which cures free radicals in the brain. In case of any type of brain disease, if taken daily If 20 basil leaves are chewed and eaten, it is very effective. Brahmi leaves: Brahmi leaves grow around our homes especially where there is soil. It is round and curved in shape. It is taken daily. Consuming it chewed on an empty stomach not only strengthens the memory but also reduces epileptic seizures. Spanish: Due to its high mercury content, the Spanish Agency for Food Safety and Nutrition recommends pregnant women not consume: A. Ripe soft cheeses, such as Brie or Camembert, B. Pâté or foie-gras, C. Calmette and Guérin bile vaccine., D. Raw sausages French: What vaccinations are required in France? A.Measles-Mumps-Rubella, B.Human papillomavirus, C.Diphtheria-Tetanus-Poliomyelitis, D.Whooping cough, E.Calmette and Guérin bile vaccine.
3.4 Localized features
In terms of symptom diagnosis, local languages retain the terminology of traditional medicine, and due to different geographical environments and living habits, the possibility that a certain symptom corresponds to different diseases is also different: for Chinese, a disease has two aspects: ”bìng” and ”zhèng”. The former is often translated as ”disease entity”. The latter, and more important one, is usually translated as ”pattern”. For example, the disease entity of a common cold might present with a pattern of wind-cold in one person, and with the pattern of wind-heat in another [^10] https://en.wikipedia.org/wiki/Traditional_Chinese_medicine#Six_Excesses.
In terms of communication terms, some languages will have some religious-related idioms at the beginning and end to improve the communication experience, such as Arabic.
4 XMedBench, Multilingual Medical Knowledge Assessment
Construction
We focus on assessing multilingual medical knowledge, select multiple-choice questions as assessment tasks, and collect common data sets with local medical characteristics as much as possible. For English, we use the MedQA-USMLE (Zhang et al., 2018), MedMCQA (Pal et al., 2022), and medical-related parts of MMLU (Hendrycks et al., 2020); for Chinese, we used the of MedQA-MCMLE (Zhang et al., 2018) and medical-related parts CMMLU (Li et al., 2023a); for Spanish, we used HEAD-QA (Vilares & Gómez-Rodríguez, 2019); for French, we used FrenMedMCQA (Labrak et al., 2023a); For Arabic and Hindi, which lack local assessments, we make a compromise by applying translated versions of MMLU [^11] Hindi: https://huggingface.co/datasets/FreedomIntelligence/MMLU_Hindi; Arabic: https://huggingface.co/datasets/FreedomIntelligence/MMLU_Arabic. Specifically, we follow Med-PaLM2 (Singhal et al., 2023) and select six subcategories in MMLU: Clinical knowledge, Medical genetics, Anatomy, Professional medicine, College biology, and College medicine. For MedQA, we choose the 4-options version. For CMMLU, we select seven subdirectories: Anatomy, Clinical knowledge, College medicine, Genetics, Nutrition, Traditional chinese medicine, and Virology.
@@ 4 XMedBench, Multilingual Medical Knowledge Assessment
Prompt User:You are a medical doctor answering real-world medical exam questions. Select one correct answer from A to D. Question: {question} Options: (A) {optiona} (B) {optionb} (C) {optionc} (D) {optiond} Assistant:The correct answer is {answer}. <special_token >
@@
Settings
We adapt 3-shot evaluation, and the specific evaluation prompt is shown in the Fig. 4. To determine the answer, we use regular matching from the answers generated by the model. For the generation strategy, we do not perform sampling and set the maximum and minimum number of generated tokens to 128 and 2. For model loading, except for the 0.5B size model which uses full precision loading, we uniformly use half precision loading. Please see the Appendix for details of the models.

### v4

@@ Abstract
Apollo: An Lightweight Multilingual Medical LLM towards Democratizing Medical AI to 6B People
[…]
@@ 1 Introduction
Figure 1: Countries covered by ApolloCorpora and relative population life expectancy
Rationale of Multilinguality in Medical LLMs The rationale of multilingual adaption in medical LLMs (Cox and Maryns, 2021; Pecina et al., 2014) could be twofold. Firstly, non-native English-speaking doctors often engage in bilingual learning through their native language and English from the outset, naturally introducing multilingual challenges in the learning process (Markó et al., 2006). Secondly, to better serve local communities, especially in countries and regions with scarce medical resources, medical aid based on local languages often achieves higher communication efficiency and acceptance (Brindley et al., 2014; Albrecht et al., 2013). Meanwhile, local medical knowledge can complement mainstream medical knowledge, fostering mutual benefits to accelerate medical development (Klayman, 1985; Yuan et al., 2016). Our pilot study in Sec. 2 also reveals that joint training of multiple languages enhances performance in the medical LLMs, indicating a beneficial complementarity among languages.
@@
The Lite LLM: Apollo The resulted multilingual medical LLMs trained by ApolloCorpora are named Apollo, to commemorate the Greek deity associated with healing, disease, the Sun, and light; this symbolizes the democratization of medical LLMs to 6 billion people, illuminating global healthcare. We explore a new domain adaption method that rewrite the pre-training corpora into QA pairs using ChatGPT (Li et al., 2023c) and adaptively sample training data, resulting in a smoother transition compared with the typical paradigm with continued pretraining and instruction tuning. Apollo ranges from 2B to 7B parameters. The advantage of the relatively-small model scale includes potential use as draft models for speculative decoding (Leviathan et al., 2023) or as proxy models for proxy-tuning (Liu et al., 2024a). In particular, we apply proxy-tuning on top of Apollo to larger general LLMs, enhancing its multilingual medical capabilities. This is achieved without the need to directly train the general model using sensitive medical corpora, thereby underscoring the practical significance of Apollo in terms of protecting the privacy of medical training data against centralized training methods.
@@
Contributions The contributions of this paper are as follows. (1) We collect and organize a high-quality multilingual medical corpora ApolloCorpora with rich local characteristics; (2) We obtain a series of SOTA multilingual LLMs Apollo at various parameter scales (especially in relatively-small sizes) ; (3) By using proxy-tuning, Apollo could significantly improve larger general LLMs without finetuning, providing a new way to mitigating the exposure of private medical training data to centralized training systems; (4) We introduce a multilingual medical evaluation XMedBench and conduct extensive benchmark of existing models.
Model English Chinese French Spanish Arabic Hindi Avg.
Base Model
Qwen-1.8B 32.91 40.07 22.12 27.43 23.71 8.82 25.84
Language Specific Models
Apollo-English 39.44 45.27 28.35 31.76 22.61 8.72 29.36
Apollo-Chinese 39.42 61.13 28.97 33.83 27.94 25.34 36.11
Apollo-French 30.94 32.71 23.81 27.00 24.54 1.74 23.46
Apollo-Spanish 33.84 43.81 27.41 35.39 28.40 23.88 32.12
Apollo-Arabic 36.40 44.27 3.74 15.73 25.90 3.03 21.85
Apollo-Hindi 25.18 3.45 18.38 19.69 1.00 25.53 15.54
Our Method
Apollo (weight average) 40.54 45.58 28.04 34.08 28.95 24.06 33.54
Table 1: The pilot study on monolingual training and multilingual training. It shows the average accuracy among datasets in each language, see details in Sec. 4.1.
@@ 2.2 The Pilot Study
Experimental settings To investigate the above question, we use a lite multilingual LLM Qwen-1.8B (Bai et al., 2023) as the LM backbone. It is argued that the findings should be agnostic to the selection of LM backbones, the selection of LLM Qwen-1.8B is due to its popularity, performance and more the importantly multilingual support. In the monolingual training setting, the LM backbone is further trained by individual language, resulted six language-specific LLM variants (i.e. , English, Chinese, French, Spanish, Arabic, and Hindi.). The training data used could be found in Sec. 3. Moreover, we average the weights of these six LLM variants and obtain a new model that does not need further training, denoted as ‘weight average’. In the multilingual training setting, we train the LM using a mixture of the corpora in these six languages.
@@
Findings. Tab. 1 highlights the effectiveness of our methods in leveraging multilingual data to enhance the performance of medical Large Language Models (LLMs). The language-specific models, each trained exclusively on data from one language, demonstrate varying degrees of improvement in their respective languages over the original LM, underscoring the value of language-specific training. However, these models show limitations outside their target languages, as seen in the relatively low scores in non-target languages, particularly in the Apollo-French-1.8B and Apollo-Hindi-1.8B models. Our method, which includes both the weight average and mmultilingual training, significantly outperforms language-specific models across all languages in terms of average performance, as shown in the last column of Tab. 1. This illustrates the substantial benefits of combining multilingual data for training medical LLMs, with marked improvements in understanding and generating medical information across a diverse set of languages.
Therefore, we conclude the finding as below.
In general, multilingual medical corpora benefits medical LLMs.
Figure 2: Overview of this work, including corpora, a benchmark and a series of medical LLMs.
@@
Prompt User:You are a medical doctor answering real-world medical exam questions. Select one correct answer from A to D. Question: {question} Options: (A) {option_a} (B) {option_b} (C) {option_c} (D) {option_d} Assistant:The correct answer is {answer}. <special_token >
@@
3 Corpora and Model Training for Apollo
3.1 ApolloCorpora: Data Collection and Cleaning
After extensive communication with doctors and medical students, we identified six high-quality medical data collections: medical books, medical encyclopedias, medical clinical guidelines, medical papers, medical examinations, and professional doctor-patient dialogues, see Tab. 2. To mimic the diverse learning experience of medical students beyond their core professional studies, we also included a wide range of medical-related content from the Internet. This approach captures the evolving nature of medical information found online. Additionally, we incorporated tasks that require mathematical reasoning and coding. This inclusion enriches the model’s skill set with critical analytical and problem-solving abilities, essential for the multifaceted demands of medical practice.
Regarding the License issue, we only screen data sets with complete open source protocols during the collection process to ensure that the open source protocols are friendly while ensuring quality. Inspired by (Cheng et al., 2023; Chen et al., 2023a), we use ChatGPT 222gpt-3.5-turbo-16k-0613 to generate questions and answers for a certain paragraph. For paragraph interception, we divide it according to the basic semantic units in the data set. Regarding the quality assurance, we rely on the help of doctors to carefully control the quality from the source of the data. The details of data collection and prepossessing are shown in App. A.
For the dimension of multilingual medical expertise, we insist on using only medical data sets entirely from local languages and do not translate any medical-related data. This is done out of the following two considerations. First, there are many related works that prove that medical translation is a very complex task that cannot be simply solved by translation software; second, the expression habits of different languages, effective drugs, and even culture and Taboo terms arising from faith need to come from the local community intact, so as to maximize communication efficiency and avoid conflicts. See localized features in App. A.3.
Data Leakage Checking The issue of data leakage is a recent focus of the academic community, which largely determines whether the results of the paper are convincing. For knowledge embedding tasks, data leakage screening with different stringency often leads to different performance. We follow Med-PaLM2 (Singhal et al., 2023) and adopt a more stringent deletion strategy. Specifically, we define a data item as leaked data if the entire question or at least 64 consecutive characters overlap with the data item. Regarding the exam exercise data source, there were 580,645 exercises before screening, and 3,041 exercises are deleted, with a screening rate of 0.52%. For other data sources, since they are not exam questions, there is no difference before and after filtering.
3.2 Apollo, the Lite Multilingual Medical LLM
We have two main starting points for training small models. First, medical equipment usually cannot call network services due to its strict privacy protection settings. For local services, the small model can achieve offline inference on the PC side, ensuring complete data localization to help improve the efficiency of medical staff; secondly, the original intention of our article is to explore a reproducible technical solution at an affordable computing cost, and promote the exploration of the field and the raising of new questions. Small models are useful for The training is very friendly for academic researchers who lack sufficient computing power.
Training models in the medical field usually involves continuing pre-training on the corpus. However, some scholars believe that although training on the original corpus gives the model domain knowledge, it greatly damages its ability to prompt question answers (Cheng et al., 2023). We consider exploring ways to rewrite the pre-training corpus into the form of question-and-answer pairs to alleviate this problem (Chen et al., 2023a). At the same time, we use priority sampling methods to achieve a smooth transition between continued pre-training and Instrcution Tuning to ensure the continuity of learning rate and data distribution transformation.
Mix Training Our dataset $D$ comprises continuing pre-training data $D_{PT}$ and instruction tuning data $D_{SFT}$. The sampling probability of each data $x\in D$ changes during training. The sampling probability of data $x$ at step $t$ during training was determined using priority sampling, defined as:
$P_{t}(x)=\frac{\pi(x)}{\sum_{y\in D-S_{t}}\pi(y)}$